I have been asked to test a piece of material that is 1/2" thick by 1" wide in a continuos length. The instruction states that specimen shall be 1/2" in thickness and 4 square inches in area. I would assume that the length needed should be 4". I have been told that it is 2" because both the top and bottom surfaces count. Is this true?
Answer by stanbon(75887)

I have been asked to test a piece of material that is 1/2" thick by 1" wide in a continuos length. The instruction states that specimen shall be 1/2" in thickness and 4 square inches in area. I would assume that the length needed should be 4". I have been told that it is 2" because both the top and bottom surfaces count. Is this true?
--------
top + bottom area = 2(1" * 2") = 4 in^2
----
Side area = 2((1/2)" * 1") + 2((1/2)" * 2") = 2[(1/2)+1) = 3 in^2
-----------------------
Cheers,
Stan H.
=================